OpenStudy (anonymous):

At what pH value does the ratio of the basic form In- to the acidic form HIn of methyl orange indicator (pKa = 4.2) have the value 5.0 : 1.0?

OpenStudy (aaronq):

Use the Henderson-Hasselbach equation: \(\sf pH=pKa+log\dfrac{[A^-]}{[HA]}\)

OpenStudy (anonymous):

pH= -log (4.2) + log (5/1)? @aaronq

OpenStudy (anonymous):

i got it wrong but it said that the answer was 4.9. could you explain?

OpenStudy (aaronq):

why did you use -log(4.2) if the pKa=4.2 ? pH= 4.2 + log (5/1) -> pH= 4.89897